# Importance sampling for Sobol’ indices estimation

### v1

@@ Abstract
We propose a new importance sampling framework for the estimation and analysis of Sobol’ indices. We show that a Sobol’ index defined under a reference input distribution can be consistently estimated from samples drawn from other sampling distributions by reweighting the estimator appropriately to account for the distribution change. We derive the optimal sampling distribution that minimizes the asymptotic variance and demonstrate its strong impact on estimation accuracy. Beyond variance reduction, the framework supports distributional sensitivity analysis via reverse importance sampling, enabling robust exploration of input distribution uncertainty with negligible additional computational cost.
@@ 1 Introduction
Complex computer models are now widely used across many scientific and engineering fields. Since they are often costly to evaluate, handling uncertainty and performing sensitivity analysis is particularly critical. In the context of global sensitivity analysis (GSA), where model inputs are treated as random variables, the aim is to understand, with as few model evaluations as possible, the global influence of each input or group of inputs on the variability of the output [8, 4].
To quantify this impact, a large variety of sensitivity indices has been proposed in the literature, each one associated with a different criterion. Some indices rely on higher-order moments [21, 24], others are based on divergence of measures between probability distributions [14, 18, 26], while others use contrast-based functionals or Cramér–von Mises-type distances [22, 33]. These indices allow for a refined view of the output distribution. Among all these tools, Sobol’ indices remain the most popular in practice. Originally introduced in [4], they are particularly well suited for scalar outputs, and can be generalized to vector-valued or functional outputs. Thanks to the Hoeffding decomposition [1], Sobol’ indices measure the contribution of each input to the output variance. As such, they are variance-based and only capture second-order effects.
In most practical applications, the model is a complex black-box code whose output distribution is unknown. Analytical formulas for Sobol’ indices are thus rarely available. As a result, these indices must be estimated, making their statistical estimation a key challenge in sensitivity analysis. Over the last decades, many estimation strategies have been proposed, which can be broadly classified into four categories: (i) Monte Carlo and quasi-Monte Carlo methods, including plain random sampling, low-discrepancy sequences, and nested designs [30, 20]; (ii) spectral and expansion techniques, such as the Fourier Amplitude Sensitivity Test (FAST) [3], Random Balance Designs (RBD) [13], the Effective Algorithm for Sensitivity Indices (EASI) [17], and polynomial chaos expansions [15]; (iii) the Pick–Freeze procedure, which directly estimates conditional variances via sample-based manipulations [29, 25]; and (iv) nonparametric estimators—including nearest-neighbor and kernel-based methods [11, 32, 36, 38, 35] and the rank-based estimator built on Chatterjee’s correlation coefficient [41, 38]—that require only a single input–output dataset.
@@
Among them, the kernel estimator of [19] is efficient and achieves the asymptotic optimal variance bound for estimating the conditional second moment $\eta=\mathbb{E}[\mathbb{E}[Y\mid X]^{2}]$.However, it remains delicate to implement and computationally demanding. On the other hand, the rank-based estimator of [41] is much simpler to implement but sub-optimal and only suitable for first-order indices.
A fundamental objective in Monte-Carlo estimation is to reduce the variance of the estimator without introducing bias. One of the most widely used techniques for variance reduction is importance sampling [2] (IS), which consists of modifying the sampling distribution to obtain a more efficient estimator of the quantity of interest. The variance of the IS estimator strongly depends on the choice of the sampling distribution.
@@
In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling. More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is often the main term requiring accurate estimation. We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution. We prove the existence of an optimal sampling distribution that can lead to zero asymptotic variance.
@@
To the best of our knowledge, no prior work has established a theoretical connection between variance reduction through importance sampling and the estimation of Sobol’ indices. A notable exception is the contribution of [23], where the authors propose a computational approach sharing similarities with importance sampling: they introduce an auxiliary variable correlated with a given input and estimate conditional expectations using a quadrature scheme (Gauss–Hermite). Their main contribution lies in the simultaneous estimation of all first-order Sobol’ indices using only a single Monte-Carlo simulation, thereby reducing numerical effort compared to classical procedures of Sobol’ indices estimation. However, their method does not address variance minimization explicitly. In contrast, our contribution focuses on the classical aspect of importance sampling for variance reduction and provides a fully flexible probabilistic framework that allows estimation from alternative sampling distributions, and also offers computational benefits presented in Section 4.
@@
The paper is structured as follows, Section 2 introduces Sobol’ indices and the classical importance sampling framework. Section 3 presents our main theoretical results. Section 4 illustrates these results with numerical experiments, including variance reduction analysis and a reverse importance sampling study of distributional sensitivity on a real test case. Section 5 concludes and outlines future research directions.
@@ 2 Main notation
Consider a triplet $(Y,\mathbf{X},\mathbf{W})$ of random variables satisfying the relation $Y=f(\mathbf{X},\mathbf{W})$ for some measurable function $f$, where $Y$ is a real-valued square integrable output, $\mathbf{X}=(X_{1},\dots,X_{k})\in\mathbb{R}^{k}$ is a vector-valued input, and $\mathbf{W}$ a random vector independent from $\mathbf{X}$ with known distribution. In the context of sensitivity analysis, the so-called Sobol’ index $S_{u}$ is used to quantify the impact of a subset $u\subseteq\{1,\dots,k\}$ of the input variables on the output $Y$, and is defined as:
$S_{u}=\frac{\mathrm{Var}\left(\mathbb{E}[Y\mid\mathbf{X}_{u}]\right)}{\mathrm{Var}(Y)}=\frac{\mathbb{E}\left[\mathbb{E}[Y\mid\mathbf{X}_{u}]^{2}\right]-\mathbb{E}[Y]^{2}}{\mathrm{Var}(Y)},$ (1)
where:
  • $\mathbf{X}_{u}=(X_{j})_{j\in u}$ denotes the subvector of $\mathbf{X}$ corresponding to the index set $u$,
  • $\mathbf{X}_{\bar{u}}=(X_{j})_{j\in\bar{u}}$ is the complementary subvector, i.e., the remaining components of $\mathbf{X}$.
As explained earlier, the main challenge for inference purposes lies in handling the conditional expectation $m_{u}(\mathbf{X}_{u})=\mathbb{E}[Y\mid\mathbf{X}_{u}]$ in order to construct an estimator of
$\eta_{u}=\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]=\mathbb{E}\left[Y\cdot m_{u}(\mathbf{X}_{u})\right]$ (2)
@@
In a given-data setting where an i.i.d. sample $(\mathbf{X}^{(i)},Y^{(i)}),\;i=1,\dots,n$ is available, optimal estimators of $\eta_{u}$ are known to be asymptotically Gaussian, with asymptotic variance achieving the efficiency bound [19] :
$\sigma^{2}_{\mathrm{opt},u}(p)=4\mathbb{E}[m_{u}^{2}(X_{u})Y^{2}]-3\mathbb{E}\left[m_{u}^{4}(X_{u})\right]-\eta_{u}^{2}.$ (3)
@@
As motivated in the introduction, we focus on improving the estimation of the key quantity $\eta_{u}$, which drives the variance of Sobol’ index estimators. For both theoretical and practical reasons, it is often beneficial to modify the original distribution of the input variables. In the importance sampling framework, briefly introduced in the next subsection, we consider the following setting: the input vector $\mathbf{X}$ is initially distributed according to a reference distribution $p$, under which the components are assumed to be dependent, although independence can also be assumed for analytical convenience. We then consider the use of an alternative sampling distribution $q$, potentially introducing a dependency structure.Our goal is to investigate whether it is possible to reduce the variance of the estimator of $\eta_{u}$ by judiciously choosing a sampling distribution $q$ and using samples $\mathbf{X}\sim q$ instead of $\mathbf{X}\sim p$.
To avoid any ambiguity, we explicitly indicate the distribution of the input $\mathbf{X}$ in the expectation notation. For example, $\mathbb{E}_{p}[\cdot]$ denotes the expectation when $\mathbf{X}$ is distributed according to the density $p$. The distribution $W$ is fixed and remains unchanged throughout. The distribution of the output $Y$ is only affected through changes in the distribution of the input.
Importance sampling (IS)
@@
A classical variance-reduction technique. To estimate an expectation of the form $\mathbb{E}_{p}[f(\mathbf{X})]$, one rewrites it under a distribution $q$ as $\mathbb{E}_{q}[f(\mathbf{X})\,w(\mathbf{X})]$ where $w(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$ is the likelihood ratio. To ensure the estimator is unbiased, the support of $q$ must include the support of $f(\mathbf{x})\,p(\mathbf{x})$. Given an i.i.d sample $\mathbf{X}^{(1)},...,\mathbf{X}^{(n)}$ drawn from $q$, the corresponding estimator is
@@
This estimator achieves zero variance if and only if $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost-everywhere. This ideal distribution is not usable in practice because it depends on the unknown quantity to be estimated. However, adaptive procedures exist to approximate it, such as non-parametric methods [7] or parametric approaches like the cross-entropy method [34, 12].
It is worth noting that the effectiveness of IS [9] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large. Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values. The next section presents our main contributions by extending the classical IS framework to characterize the optimal sampling distribution $q$ for estimating $\eta_{u}$ and improving Sobol’ indices estimation.
@@ 3.1 Input distribution change via importance sampling
We consider a modification of the input distribution, applied exclusively to the random vector $\mathbf{X}\in\mathbb{R}^{k}$. Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$, for any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\textbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x})$ its marginal density and introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
$q(\mathbf{x})=q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}),$ (5)
where $q_{u}(\mathbf{x}_{u})$ and $q_{\bar{u}}(\cdot\mid\mathbf{x}_{u})$ denote the marginal and conditional densities, respectively. The corresponding importance weights are then given by:
$w(\mathbf{x})=\frac{p(\mathbf{x})}{q(\mathbf{x})}=\frac{p_{u}(\mathbf{x}_{u})\cdot p_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}{q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}=w_{u}(\mathbf{x}_{u})\cdot w_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}).$ (6)
Lemma
Let $q=q_{u}\cdot q_{\bar{u}}$ be any importance sampling distribution on the input $\mathbf{X}$, define
$Z_{u}:=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot Y.$ (7)
@@
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}).$ (8)
@@
$\eta_{u}=\mathbb{E}_{p}\left[(\mathbb{E}_{p}[Y\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{2}\right]=\mathbb{E}_{q}\left[Z_{u}\cdot\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right].$ (9)
Proof
See Appendix
This identity provides a reformulation of $\eta_{u}$ under the sampling distribution $q$ and can be consistently estimated using the auxiliary variable $Z_{u}$. We will show next that a suitable choice of $q$ leads to a reduced variance of the corresponding estimator.

### v2

@@ Abstract
We propose a new importance sampling framework for the estimation and analysis of Sobol’ indices. We focus on the estimation of the conditional second-moment quantity underlying these indices, which is the most challenging term to estimate. We show that this quantity, originally defined under a reference input distribution, can be estimated from samples drawn under auxiliary distributions by reweighting the model outputs. We derive the optimal sampling distribution that minimises the asymptotic variance of efficient estimators and demonstrate its impact on estimation. Beyond variance reduction, the framework also supports distributional sensitivity analysis through reverse importance sampling.
@@ 1 Introduction
Complex computer models are now widely used across many scientific and engineering fields. Since they are often costly to evaluate, handling uncertainty and performing sensitivity analysis is particularly critical. In the context of global sensitivity analysis (GSA), where model inputs $\mathbf{X}=(X_{1},\dots,X_{k})$ are treated as random variables, the aim is to understand, with as few model evaluations as possible, the global influence of each input or group of inputs on the variability of the model output $Y$ [41, 42].
To quantify this impact, a large variety of sensitivity indices have been proposed in the literature, each one associated with a different criterion. Some indices rely on higher-order moments [34, 36], others are based on discrepancy measures between probability distributions [4, 5, 10], while others use contrast-based functionals or Cramér–von Mises distance [20, 24]. These indices allow for a refined view of the output distribution. Among all these tools, Sobol’ indices remain the most popular in practice. Originally introduced in [42], they are particularly well suited for scalar outputs, and can be generalised to vector-valued or functional outputs. Thanks to the Hoeffding decomposition [27], Sobol’ indices measure the contribution of each group of inputs to the output variance. The first-order Sobol’ indices quantify the main effects of each input, while the total-order Sobol’ indices also account for both main and interaction effects. Hence, they are variance-based measures that capture the contribution of inputs to the overall variability of the model output.
In most practical applications, the model is a complex black-box code whose output distribution is unknown. Analytical formulas for Sobol’ indices are thus rarely available. As a result, these indices must be estimated, making their statistical estimation a key challenge in sensitivity analysis, especially for the conditional second moment $\eta=\mathbb{E}[\mathbb{E}[Y\mid X]^{2}]$. Over the last decades, many estimation strategies have been proposed and can be broadly classified into four categories: (i) Monte Carlo and quasi-Monte Carlo methods, including plain random sampling, low-discrepancy sequences, and nested designs [25, 35]; (ii) spectral and expansion techniques, such as the Fourier Amplitude Sensitivity Test (FAST) [9], Random Balance Designs (RBD) [47], the Effective Algorithm for Sensitivity Indices (EASI) [39], and polynomial chaos expansions [45]; (iii) the Pick–Freeze procedure, which directly estimates conditional variances via sample-based manipulations [23, 28]; and (iv) nonparametric estimators including nearest-neighbor [17, 16, 7] and kernel-based methods [11, 21, 12], and the rank-based estimator [30, 22] built on Chatterjee’s correlation coefficient [8] which share the practical advantage of being applicable in the so-called given-data setting, where only a fixed input–output sample set is available and direct access to the black-box model is not possible.
@@
With the increasing complexity of models and growing computational costs, the Pick–Freeze setting is becoming less practical in many applications, as it requires two evaluations of the model for each realization of $X$, which is often prohibitive. When the context allows it, using this particular design can considerably simplify estimation, but the number of code calls often makes it prohibitive. By contrast, given-data methods are increasingly favored because they operate from a provided input–output sample only. This naturally raises the question of efficiency limits in the given-data setting.
Among them, several estimators achieve the asymptotic efficiency lower bound. The expression of the efficiency bound for $\eta$ was first derived in [18] for a continuous input $X$, from the efficient influence function associated with the parameter. The same work also proposed an efficient estimator of a truncated version of $\eta$. Since then, several methods have been shown to be asymptotically efficient in the given-data setting. For first-order indices, the kernel estimator introduced in [11] achieves the asymptotic optimal variance bound for estimating $\eta$ (see Theorems 3.4 and 3.5). This result was initially established for first-order Sobol’ indices and later extended to higher-order ones under suitable regularity conditions [12] (see Theorems 4.1 and 4.2). These estimators, however, remain difficult to implement in practice. In contrast, the rank-based estimator of [22], built on order statistics, is much simpler but sub-optimal and restricted to first-order indices, although it has been further generalized to attain the efficiency bound in [31] (Proposition 3.3). The question of efficiency bounds has been thoroughly discussed in [30]. In this work, we mainly focus on these efficient estimators.
Even though these estimators are already asymptotically optimal under the reference input distribution, their accuracy may degrade when the sample size is small or when the model response is highly nonlinear. In such cases, a large fraction of the simulation effort may be wasted on regions of the input space that contribute little to the Sobol’ quantity of interest, while the most informative regions are poorly explored. When the black-box model is still available, a natural way to improve efficiency is to modify the sampling scheme so that evaluations are more concentrated on informative regions. This motivates the use of importance sampling (IS) [29], one of the most classical and effective approaches: it modifies the sampling distribution to obtain a more accurate estimator of the target quantity. The variance achieved by IS strongly depends on how this sampling distribution is chosen.
@@
In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling. More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is the most challenging term to estimate accurately. We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution, which makes it possible to characterize optimal sampling distribution for efficient estimators, potentially achieving zero variance under appropriate conditions.
@@
To the best of our knowledge, no prior work has established a theoretical connection between variance reduction through importance sampling and the estimation of Sobol’ indices. Two related contributions deserve to be mentioned. First, [2] propose a computational approach sharing similarities with importance sampling: they introduce an auxiliary variable correlated with a given input and estimate conditional expectations using a Gauss–Hermite quadrature scheme. Their main contribution lies in the simultaneous estimation of all first-order Sobol’ indices using only a single Monte Carlo simulation, thereby reducing numerical effort compared to classical procedures. However, their method does not address variance minimisation explicitly. Second, [6] introduce a weighting method for estimating Sobol’ total effects under constrained input spaces, where an importance-type density ratio is used to correct Jansen’s estimator in non-Cartesian domains. In contrast, our contribution focuses on the classical aspect of importance sampling for variance reduction and provides a fully flexible probabilistic framework that also extends naturally to distributional sensitivity analysis.
Beyond variance reduction, the proposed importance sampling framework can also be exploited to assess the robustness of Sobol’ indices with respect to changes in the input distributions. This corresponds to the concept of reverse importance sampling [3, 1, 32], which exploits the fact that a single dataset can be reused to evaluate sensitivity measures under alternative probabilistic scenarios. Note that this should not be confused with the reverse importance sampling estimator (or harmonic mean estimator) used in Bayesian model evidence estimation [33]. A post-processing importance sampling approach was also proposed in [46], where existing Monte Carlo samples are reweighted to estimate global sensitivity indices. Their method relies on a kernel-like weighting scheme to approximate conditional moments.
@@
The paper is structured as follows. Section 2 introduces Sobol’ indices and the classical importance sampling framework. Section 3 presents our main theoretical results. Section 4 illustrates these results with numerical experiments, including variance reduction analysis and a reverse importance sampling study of distributional sensitivity on a real-life test case. Section 5 concludes and outlines future research directions.
@@ 2 Main notation
Consider a triplet $(Y,\mathbf{X},\mathbf{W})$ of random variables satisfying the relation $Y=f(\mathbf{X},\mathbf{W})$ for some measurable function $f$, where $Y$ is a real-valued square-integrable output, $\mathbf{X}=(X_{1},\dots,X_{k})\in\mathbb{R}^{k}$ is a vector-valued input, and $\mathbf{W}$ is a random vector independent of $\mathbf{X}$. In the context of sensitivity analysis, the so-called Sobol’ index $S_{u}$ is used to quantify the impact of a subset $u\subseteq\{1,\dots,k\}$ of the input variables on the output $Y$, and is defined as:
$S_{u}=\frac{\mathrm{Var}\!\left(m(\mathbf{X}_{u})\right)}{\mathrm{Var}(Y)},$ (1)
where $m(\mathbf{X}_{u})=\mathbb{E}[Y\mid\mathbf{X}_{u}=\mathbf{x}_{u}]$, with $\mathbf{X}_{u}=(X_{j})_{j\in u}$ the subvector of inputs corresponding to the index set $u$, and $\mathbf{X}_{-u}=(X_{j})_{j\notin u}$ its complementary subvector containing the remaining components of $\mathbf{X}$. One can develop the variance term as $\mathrm{Var}(m(\mathbf{X}_{u}))=\mathbb{E}[m^{2}(\mathbf{X}_{u})]-(\mathbb{E}[Y])^{2}$, and both expectations and variances can in practice be estimated through classical Monte Carlo estimators based on independent samples of $(Y,\mathbf{X})$. Hence, as recalled above, the main challenge for inference purposes lies in handling the conditional expectation $m(\mathbf{X}_{u})$ itself in order to construct an estimator of
$\eta_{u}=\mathbb{E}\!\left[m^{2}(\mathbf{X}_{u})\right]=\mathbb{E}\!\left[m(\mathbf{X}_{u})\,\mathbb{E}[Y\mid\mathbf{X}_{u}]\right]=\mathbb{E}[\,\mathbb{E}[Y\,m(\mathbf{X}_{u})\mid\mathbf{X}_{u}]\,]=\mathbb{E}\!\left[Y\,m(\mathbf{X}_{u})\right],$ (2)
where the second equality follows from the law of total expectation.
@@
For an i.i.d. sample $(\mathbf{X}^{(i)},Y^{(i)}),\;i=1,\dots,n$, efficient estimators of $\eta_{u}$ are asymptotically Gaussian, with variance achieving the lower bound [18]. Writing $\nu(\mathbf{X}_{u})=\mathrm{Var}(Y\mid\mathbf{X}_{u})$, the optimal asymptotic variance is given by:
$\sigma^{2}_{\mathrm{opt},u}=4\,\mathbb{E}\!\left[m^{2}(\mathbf{X}_{u})\,\nu(\mathbf{X}_{u})\right]+\mathrm{Var}\!\left(m^{2}(\mathbf{X}_{u})\right).$ (3)
@@
As motivated in the introduction, we focus on improving the estimation of the key quantity $\eta_{u}$. For both theoretical and practical reasons, it is often beneficial to modify the original distribution of the input variables. Our goal is to investigate whether it is possible to reduce the variance of the estimator of $\eta_{u}$ by judiciously choosing a sampling distribution $q$.
In the importance sampling framework, briefly recalled below, the input vector $\mathbf{X}$ is initially drawn from a reference distribution $p$, under which the components may be dependent, although independence can be assumed for analytical convenience. We then introduce an alternative sampling distribution $q$, possibly with a different dependency structure. To avoid ambiguity, expectations are indexed by the underlying distribution: for instance, $\mathbb{E}_{p}[\cdot]$ denotes expectation under $p$. The distribution of $\mathbf{W}$ is fixed and remains unchanged, and the output $Y$ is affected only through changes in the distribution of $\mathbf{X}$.
Importance sampling
@@
This classical variance-reduction method estimates expectations of the form $\mathbb{E}_{p}[f(\mathbf{X})]$ by rewriting them under another distribution $q$ as $\mathbb{E}_{q}[f(\mathbf{X})\,w(\mathbf{X})]$, where $w(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$ denotes the likelihood ratio. To ensure unbiasedness, the support of $q$ must include that of $f(\mathbf{x})\,p(\mathbf{x})$. Given an i.i.d. sample $\mathbf{X}^{(1)},\dots,\mathbf{X}^{(n)}$ drawn from $q$, the corresponding estimator is
@@
This estimator achieves zero variance if and only if $f$ is nonnegative and $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost everywhere. This ideal distribution cannot be used in practice because it depends on the unknown quantity to be estimated. However, adaptive procedures exist to approximate it, such as nonparametric methods [49] or parametric approaches like the cross-entropy method [38, 14]. The efficiency of IS [37] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large. Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values.
The next section extends the classical IS framework to characterize the optimal $q$ for estimating $\eta_{u}$ and improving the estimation of Sobol’ indices.
@@ 3.1 Input distribution change via importance sampling
We introduce the importance sampling framework by allowing a modification of the input distribution of the random vector $\mathbf{X}\in\mathbb{R}^{k}$, while keeping the noise variable $\mathbf{W}$ fixed. Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$. For any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\mathbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x}_{u})$ its marginal density. We introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
$q(\mathbf{x})=q_{u}(\mathbf{x}_{u})\,q_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}),$ (5)
where $q_{u}(\mathbf{x}_{u})$ and $q_{-u\mid u}(\cdot\mid\mathbf{x}_{u})$ denote the marginal and conditional components of $q$. The corresponding importance weights are then given by
$w(\mathbf{x})=\frac{p(\mathbf{x})}{q(\mathbf{x})}=\frac{p_{u}(\mathbf{x}_{u})\,p_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})}{q_{u}(\mathbf{x}_{u})\,q_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u})}=w_{u}(\mathbf{x}_{u})\,w_{-u\mid u}(\mathbf{x}_{-u}\mid\mathbf{x}_{u}).$ (6)
Lemma 3.1
Let $q=q_{u}\,q_{-u\mid u}$ be any importance sampling distribution on the input $\mathbf{X}$, and define
$Z_{u}:=\sqrt{w_{u}(\mathbf{X}_{u})}\,w_{-u\mid u}(\mathbf{X}_{-u}\mid\mathbf{X}_{u})\,Y.$ (7)
@@
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]=\sqrt{w_{u}(\mathbf{X}_{u})}\,m(\mathbf{X}_{u}).$ (8)
@@
$\eta_{u}=\mathbb{E}_{p}\!\left[(\mathbb{E}_{p}[Y\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\!\left[(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\!\left[Z_{u}\,\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right].$ (9)
The proof is given in the Appendix.
In the following, this identity is used to express $\eta_{u}$ under $q$ and to show that an appropriate choice of $q$ can reduce the variance of its estimator.